Help me solve this physics homework:
in sense we all have kinetic energy,even when we are standing still.the earth with a radius of 6.37*10 power 6 m,rotates about its axis once a day.ignoring the earths rotation about the sun,what is the K.E of a 50Kg man standing on the surface of the earth.
1
Expert's answer
2013-03-05T09:23:28-0500
For the kinetic energy we have the following formula:
T = m*v^2/2
T - kinetic energy
m - mass
v - velocity
To determine the velocity we apply this one (as the Earth is rotating):
v = 2*pi*R/T
R - radius of the Earth
T - period of rotation ( 1 day )
pi = 3.14
v = 2*3.14*6.37*10^6 m / (24*3600 sec ) = 463 m/s
T = m*v^2/2 = 50*463^2/2 J = 5359225 J = 5.36 MJ